### v2

# Interpretable Company Similarity with Sparse Autoencoders

Abstract
Determining company similarity is a vital task in finance, underpinning hedging, risk management, portfolio diversification, and more. Practitioners often rely on sector and industry classifications to gauge similarity, such as SIC-codes and GICS-codes – the former being used by the U.S. Securities and Exchange Commission (SEC), and the latter widely used by the investment community. Since these classifications can lack granularity and often need to be updated, using clusters of embeddings of company descriptions has been proposed as a potential alternative, but the lack of interpretability in token embeddings poses a significant barrier to adoption in high-stakes contexts. Sparse Autoencoders (SAEs) have shown promise in enhancing the interpretability of Large Language Models (LLMs) by decomposing LLM activations into interpretable features. We apply SAEs to company descriptions, obtaining meaningful clusters of equities in the process. We benchmark SAE features against SIC-codes, Industry codes, and Embeddings. Our results demonstrate that SAE features not only replicate but often surpass sector classifications and embeddings in capturing fundamental company characteristics. This is evidenced by their superior performance in correlating monthly returns – a proxy for similarity – and generating higher Sharpe ratio co-integration strategies, which underscores deeper fundamental similarities among companies.
LLM
Large Language Model
LLMs
Large Language Models
LMs
Language Models
LM
Language Model
MI
Mechanistic Interpretability
SAE
Sparse Autoencoder
SAEs
Sparse Autoencoders
Interpretable Company Similarity with Sparse Autoencoders
Marco Molinari*, 1and Victor Shao*, 1and Vladimir Tregubiak*, 1 Abhimanyu Pandey1  and Mateusz Mikolajczak1  and Sebastião Kuznetsov Ryder Torres Pereira1 1 LSE.AI, London School of Economics and Political Science https://lseai.org Correspondence: m.molinari1@lse.ac.uk *Equal Contribution
1 Introduction
Accurately assessing the similarity of companies is an integral task in finance, key to risk management and portfolio diversification (Delphini et al., 2019; Katselas et al., 2017). Hedging, a practice that relies on converse investments in related assets, is a prominent example of a financial strategy that requires a detailed understanding of the similarity between two companies.
Traditional modes of company comparisons rely on relative returns and discrete classifications, e.g. SIC codes or the Global Industry Classification System (GICS), which categorizes companies into 11 sectors and 163 sub-industries (MSCI, 2020). Relying on return spreads can be effective but is not foolproof, as market volatility, economic shifts, business fundamental changes, and temporal factors can disrupt them (Loretan and English, 2000).
Additionally, systems like GICS are limited, as the restricted granularity of a discrete classification system limits dynamic interpretations of companies’ operations, as it falls short in accounting for the duality of certain companies, particularly in emerging industries (Winton, 2018).
This is particularly important for pairs trading, a market-neutral strategy based on mean-reverting return spreads (Ehrman, 2012). Employing a pairs trading strategy with fundamentally similar companies whose returns are co-integrated could lessen the risk of deviation from historical trends (Raghava and Bharadwaj, 2014).
Clustering embeddings of company descriptions has been proposed as a measure of similarity (Vamvourellis et al., 2023), but token embeddings are not interpretable, and this uncertainty is undesirable in the financial sector.
SAEs have the potential to provide an efficient measure of company similarity by decomposing large amounts of financial data into interpretable features (Chen et al., 2020). Moreover, SAEs have recently been applied to LLMs resulting in interpretable decompositions of neural activations (Huben et al., 2024). This motivates their application to textual company descriptions. We are the first, to the best of our knowledge, to compute company similarity using SAEs and SEC filings.
1.1 Sparse AutoEncoders
The linear representation hypothesis posits that LLMs linearly represent concepts in neuron activations (Park et al., 2024). However, neuron activations are notoriously superpositioned (Elhage et al., 2022), SAEs enhance the interpretability of LLMs by writing neuron activations as a linear combination of sparse features (Bricken et al., 2023). This reduces the superposition and restores the interpretability (Huben et al., 2024). SAEs have recently been applied both in mechanistic interpretability (Nanda et al., 2023; Conmy et al., 2023; Marks et al., 2024) of LLMs, and in deep learning more broadly (Chen and Guo, 2023). SAEs have been scaled to medium size and large Language Models (LMs), such as GPT4 (Templeton et al., 2024; Gao et al., 2024).
The Sparse Autoencoder (SAE) learns a reconstruction $\hat{\mathbf{x}}$ as a sparse linear combination of features $\mathbf{y}_{i}\in\mathbb{R}^{d_{\text{s}}}$ for a given input activation $\mathbf{x}\in\mathbb{R}^{d_{\text{m}}}$ with $d_{m}$ as the Large Language Model (LLM)’s hidden size and:
$d_{\text{s}}=k\,d_{\text{m}},\quad\text{with }k\in\{2^{n}\mid n\in\mathbb{N}_{+}\}.$ (1)
The decoder element of the SAE is given as:
$(\hat{\mathbf{x}}\circ\mathbf{f})(\mathbf{x})=\mathbf{b}_{d}+\mathbf{W}_{d}\mathbf{f}(\mathbf{x})$ (2)
where $\mathbf{b}_{d}\in\mathbb{R}^{d_{\text{m}}}$ is the bias term of the decoder, and $\mathbf{W}_{d}$ is the decoder matrix with columns $\mathbf{v}_{i}\in\mathbb{R}^{d_{\text{m}}}$, and $\mathbf{f}(\mathbf{x})$ denotes the feature activations, which are described by:
$\mathbf{f}(\mathbf{x})=\text{TopK}(\mathbf{W}_{e}(\mathbf{x}-\mathbf{b}_{d})+\mathbf{b}_{e})$ (3)
where $\mathbf{b}_{e}\in\mathbb{R}^{d_{\text{s}}}$ is the bias term of the encoder, and $\mathbf{W}_{e}$ is the decoder matrix with columns $\mathbf{w}_{i}\in\mathbb{R}^{d_{\text{s}}}$. The loss function, which also corresponds to the output’s mean-squared error (MSE) is defined by:
$\mathcal{L}_{\text{}}=\norm{\mathbf{x}-\hat{\mathbf{x}}^{2}_{2}}\$ (4)
The choice of a TopK function rather than ReLU or L1 Regularization autoencoders is due to TopK having been empirically shown to outperform the latter in balancing sparsity and reconstruction accuracy (Gao et al., 2024).
1.2 Our contributions
  • We apply an open source SAE (EleutherAI, 2024) to Llama 3 8B (Dubey et al., 2024), and extract interpretable features from SEC filings.
  • We release a dataset containing company descriptions, extracted features, and returns, to support further research.
  • We train a Siamese Network to obtain expected correlations from company descriptions, and construct a minimum spanning tree of stocks where correlated stocks are connected. We then perform hierarchical clustering with results comparable to, or better, than Embeddings, SIC-codes, and Industry categorizations.
  • We demonstrate that a combined model, integrating SAE features with SIC-codes and Industry categorizations, significantly outperforms standalone approaches in explaining pairwise correlations, showing that our method can enhance industry classifications.
  • We validate the applicability of our features via co-integration pairs trading strategies, and show that a strategy based on sparse features leads to a higher Sharpe ratio, demonstrating the potential for wider uses within finance.
2 Methodology
2.1 Dataset
Publicly listed companies in the U.S. submit annual financial performance reports to the Securities and Exchange Commission (SEC) which include information on a company’s operations, such as product specifics, subsidiaries, competition, and other financial details(SEC, 2023) . We consider 220,275 annual reports from 1993 to 2020, ignoring any de-lists, accompanied by related meta-data on Company Name, Year, SIC-code, and CIK number (a unique SEC corporation identifier). CIK numbers are mapped to their corresponding publicly traded ticker symbol, from which the monthly logged returns are retrieved via Yahoo Finance (2024). Due to a lack of publicly available mappings between CIK codes and GICS classifications, we utilize: (1) SIC-codes, and (2) the industry/major-division categorization [^1] The first 3 digits of the SIC code splits companies into 12 industry/major-divisions, referred to hereafter as BISC (Broader Industry Sector Code). (U.S. Occupational Safety and Health Administration, 2001). We remove entries with missing or very short: company descriptions, ticker information, or monthly returns. Next, company descriptions are tokenized using Meta’s Llama 3 8B Tokenizer (Dubey et al., 2024). We further refine the dataset by retaining only annual reports that have token counts within the context window, and are consistently available for at least five years, leaving us with a dataset of 27,722 reports. In our analysis, we ignore pre-1996 as the sample size is too small.
We employ forward testing in our experiments: when any training is done on year $t$, we report and use the predicted results on $t+1$. Our dataset analysis finds that only 0.24% of subsequent company descriptions are identical, with 0.54% within a normalized Levenshtein distance of 0.01, and 12.34% within 0.05. Forward-testing is more robust than out-of-sample testing (Pardo, 1992), and since the descriptions are not repeated across years, the model remains generalized.
2.2 Feature summing
In this work, we face the challenge of comparing sparse feature sequences of arbitrary lengths, where best practices are not well-established, though max-pooling has been proposed as a baseline for feature aggregation (Bricken et al., 2024). However, motivated by the specific demands of financial sequence modeling, we propose an alternative, employing sparse feature summing across tokens. This method provides a magnitude-scaled count of how frequently a feature appears within a sequence, reflecting both the number of tokens a feature is active for and its intensity (Lan et al., 2024).
Our approach is inspired by analogous methodologies in literature. For example, Loughran et al. (2009) highlight the value of summing word counts in financial text analysis to derive domain insights.
We sum sparse features, across tokens, from an SAE (EleutherAI, 2024) applied at the layer occurring 90% of the way through the model. At this layer, we capture relevant features from preceding layers via the skip connection (Vaswani et al., 2017), but not the logic-related features that tend to occur at the very last layers (Ghilardi et al., 2024).
The skip connection ensures that a single SAE captures the entire residual stream (Longon, 2024), inherently including information from all preceding layers, thus ensuring that the summed sparse features represent a comprehensive aggregation of the descriptions’s representation in the model.
We analyze summed sparse features, and observe an interesting exponential decay pattern in feature activation frequencies, where half the features never activate, a fourth activate once, an eighth activate twice, and so forth (Figure 1).
Figure 1 highlights the sparsity of LLM latent features – even when these are summed across thousands of tokens – motivating feature summing as an approach. In this context, before summing, a single active feature has, on average, a value of $\approx{0.7}$ (the first bulge).
This method also addresses a limitation in using embeddings as opposed to sparse features (Vamvourellis et al., 2023), which requires equal-length sequences for comparison. By focusing on cumulative feature occurrences, summed sparse features enable comparisons between sequences of arbitrary lengths, offering greater flexibility for analyzing variable-length financial datasets.
Figure 1: Distribution of summed feature activations for the whole dataset.
2.3 Models
2.3.1 Embedders
As a baseline, we replicate Vamvourellis et al. (2023)’s embedding methodology and obtain embeddings for company descriptions in our dataset. In particular, we use:
  1. 1.
    BERT: Given the maximum token length of 512, we employ chunking by splitting company descriptions into overlapping chunks of 512 tokens, extracting CLS embeddings for each chunk, and averaging these embeddings to generate a single 1536-token representation, following (Vamvourellis et al., 2023).
  2. 2.
    Sentence-BERT (SBERT): Building on BERT, SBERT improves latency substantially (Reimers and Gurevych, 2019) and encodes meaning on the more abstract sentence level. We again employ chunking to 1536 tokens.
  3. 3.
    PaLM-gecko: Pathways Language Model (PaLM) (Chowdhery et al., 2022).
2.4 Siamese Networks
Given the comparative nature of our use case, we employ Siamese neural networks (Bromley et al., 1993) to validate the meaningfulness of the extracted features. Specifically, the Siamese network is trained to predict the correlation $p_{i,j}$ between two companies $i$ and $j$  [^2] Monthly returns are used over higher-frequency returns to reduce noise from idiosyncratic events (Huang et al., 2009)..
The Siamese network uses similarity information to position correlated companies closer together in the feature space. It employs two identical sub-networks with shared weights and parameters (Koch et al., 2015). Features extracted by the SAE from companies $A$ and $B$, denoted as $f_{1}$ and $f_{2}$, are passed through two dense layers, each followed by a ReLU activation, to produce encoded representations $z_{1}$ and $z_{2}$. Cosine similarity of $z_{1}$ and $z_{2}$ is then used in the MSE loss function for network optimization [^3] See Appendix E. We train a new Siamese network for each year, using data from $t$ (1993–2019) to predict $t+1$ (1994–2020), iterating year by year.
2.5 Clustering
We benchmark our sparse features against embeddings and SIC/BISC-codes, where each SIC/BISC-code is its own cluster.
[…]
$\text{MeanCorr}(G_{k})=\frac{1}{|G_{k}|}\sum_{C_{i}\in G_{k}}\frac{1}{|C_{i}|}\sum_{(x,y)\in C_{i}}\rho(x,y)$
where $\rho(x,y)$ is the Pearson correlation of logged monthly returns for companies $x$ and $y$. This metric quantifies stock return coherence within clusters, evaluating each clustering method’s effectiveness.
2.5.1 Clustering Embeddings
Each of the embedders we employ is used to define a group/set-of-clusters ($G_{\text{embedder}}$): (a) $G_{\text{BERT}}$; (b) $G_{\text{SBERT}}$; (c) $G_{\text{PaLM-gecko}}$ following Vamvourellis et al. (2023). We first reduce the dimensionality of the embeddings using UMAP (McInnes et al., 2020), then generate clusters through Spectral Clustering. See Appendix A for methodology details.
2.5.2 Clustering Sparse Features
Sparse features lack the locality and smoothness of embeddings (Kiros et al., 2015; Bischke et al., 2019) to define reliable similarity metrics. Moreover, half of the features never activate (See Figure 1) and hence are unrelated to finance, but are present due to the SAE being trained on a generic dataset (Weber et al., 2024). Thus, we cannot directly cluster the sparse features.
To cluster the sparse features, we adopt the graph-theoretic framework of Bonanno et al. (2004), employing Minimum Spanning Trees (MSTs) to extract hierarchical structures from financial data. A fully connected graph is constructed with edge weights representing a particular distance metric (see metrics 1, 2 below). The MST encodes a subdominant ultrametric, with ultrametric distance defined by the maximum edge weight on the unique path between two nodes [^4] To enforce the ultrametric property, we employ single-linkage hierarchical clustering, which groups nodes by iteratively merging the pair of clusters with the smallest maximum distance between any two points. This process inherently satisfies the ultrametric inequality $(d_{ij}\leq\max(d_{ik},d_{kj}))$ by construction.. We remove edges above a specified weight level [^5] See Appendix B on edge-weight cutoff optimisation, generating clusters directly from the MST. This eliminates the need for additional clustering steps, ensuring stable and interpretable results consistent with Bonanno et al. (2004).
Metric 1: Siamese-derived Metric: This cluster group is constructed using a siamese-derived ultrametric, and will be called $G_{\text{C-TS}}$. First we obtain predicted correlations from the Siamese (SN) model. Then, following (Bonanno et al., 2004) we use these to define a distance metric between stock pairs, where $\Delta t$ is the time horizon:
$d_{i,j}(\Delta t)=\sqrt{2(1-\rho_{i,j}(\Delta t))}.$
The distance matrix $\mathbf{D}(\Delta t)$ [^6] With this choice, $d_{i,j}(\Delta t)$ fulfills the three axioms of a metric: (i) $d_{i,j}(\Delta t)=0$ iff $i=j$; (ii) $d_{i,j}(\Delta t)=d_{j,i}(\Delta t)$; (iii) $d_{i,j}(\Delta t)\leq d_{i,k}(\Delta t)+d_{k,j}(\Delta t)$. is then used to determine the MST connecting the $n$ stocks.
Metric 2: Manhattan Metric: This group, $G_{\text{C-TM}}$, is based on a second metric, $M_{i,j}$, derived from PCA-transformed sparse features, where 131k-dimensional vectors are reduced to 4,000 dimensions [^7] We fit PCA globally across 1996–2020 for consistent eigenvectors.. For each pair of companies $i$ and $j$, $M_{i,j}$ is computed as:
$M_{i,j}=\sum_{k=1}^{4000}\left|\text{PCA}_{k}(\mathbf{f}_{i})-\text{PCA}_{k}(\mathbf{f}_{j})\right|$
where $M_{i,j}$ serves as a metric888 See Appendix C for the full proof of dissimilarity (Kotsiantis et al., 2001).
2.6 Pairs Trading
Pairs trading assumes correlated stock returns, with traditional strategies often focusing on relative performance diagnostics such as price ratio spreads (Fallahpour et al., 2016).
The dataset is divided into an in-sample period (Jan 2002–Dec 2013) and an out-of-sample period (Jan 2014–Dec 2020), with clusters $G_{k}$ ($G_{\text{C-T1}}$, SIC-codes, BISC, and their joined variations). The pairs trading strategy involves the following steps:
  1. 1.
    Pre-selection: For each cluster $C_{i}\in G_{k}$, stock pairs are filtered if the Pearson correlation of their monthly logged returns exceeds 0.9 during the in-sample period.
  2. 2.
    Co-integration Testing: Using historical adjusted close data (Jan 2002–Dec 2013) for the pre-selected pairs while standardized for corporate action, an Engle-Granger co-integration test is conducted with the Augmented Dickey-Fuller (ADF) statistic to determine whether the residual spread is stationary. Pairs with a p-value below 0.05 are considered co-integrated.
  3. 3.
    Ranking and Selection: Co-integrated pairs are ranked by their ADF test p-values [^9] We choose pairs with lower p-values, as they indicate stronger stationarity of the spread, reflecting a stronger co-integration relationship, with the top 80 pairs retained for out-of-sample evaluation (see Table 1). The effectiveness of co-integration within each cluster group $Gk$ is assessed by evaluating the Sharpe ratio of the entire portfolio. The Sharpe ratio quantifies risk-adjusted returns, measuring excess return per unit of risk. We prioritized this over absolute PnL, as higher Sharpe ratios enable leverage to amplify returns, provided practical constraints such as transaction costs and risk limits are effectively managed (Guasoni and Mayerhofer, 2018; Peters, 2011).
See Appendix D for Trading Logic Detail.
3 Results
3.1 Clustering Results
Figure 2: Mean correlation of C-TS (Siamese) Cluster vs Benchmarks between 1996-2020. Joint model significantly outperforming benchmarks, suggesting sparse features and SIC-codes capture complementary, non-overlapping information.
[…]
Clustering Group $\mathbf{(G_{k})}$ $\text{MeanCorr}\mathbf{(G_{k})}$ (1996-2020) Sharpe Ratio
$\mathbf{G_{\text{C-TS}}}$ $\mathbf{0.246}$ $\mathbf{12.97}$
$\mathbf{G_{\text{C-TM}}}$ $\mathbf{0.266}$ $\mathbf{15.84}$
$G_{\text{BERT}}$ 0.198 12.68
$G_{\text{SBERT}}$ 0.219 9.34
$G_{\text{PaLM-gecko}}$ 0.219 9.97
$\mathbf{G_{\text{C-TS $\cap$ SIC }}}$ $\mathbf{0.454}$mathbf{-}-
$\mathbf{G_{\text{C-TS $\cap$ BISC}}}$ $\mathbf{0.327}$ $\mathbf{13.88}$
$\mathbf{G_{\text{C-TM $\cap$ SIC}}}$ $\mathbf{0.416}$mathbf{-}-
$\mathbf{G_{\text{C-TM $\cap$ BISC}}}$ $\mathbf{0.412}$mathbf{-}-
$G_{\text{BERT $\cap$ SIC}}$mathbf{-}- 12.11
[…]
$G_{\text{SBERT $\cap$ BISC}}$ 0.292 14.66
[…]
SIC 0.231 10.73
BISC 0.187 13.85
[…]
3.2 Pairs Trading Results
During the live trading simulation, cumulative PnL (profits) (Smith and Doe, 2021) and Sharpe ratios (risk-adjusted profits) were recorded for evaluation. Results are reported in Figure 3.
Figure 3: Cumulative PnL and Sharpe ratios comparison between $G_{\text{C-TM}}$ Cluster and Benchmarks: $G_{\text{C-TM}}$ significantly outperforms in terms of Sharpe
While the co-integration test ensures statistical validity, the strategy’s profitability relies on the underlying structural fundamental similarities between companies. Our clustering approach outperforms SIC & BISC groups in Sharpe ratio, suggesting our method captures more fundamental aspects of company similarity (Hong and Hwang, 2021).
3.3 Interpretability
Sparse features are understood to be interpretable (Huben et al., 2024), and our core contribution is showing that it is possible to match or improve upon existing methodologies (U.S. Occupational Safety and Health Administration, 2001; Vamvourellis et al., 2023) while imposing an interpretability constraint.
In practice, we use interpretability to ascertain the economic rationale underlying the most co-integrated stock pairs within each cluster $C_{i}$ in $G_{\text{C-TM}}$. This allows us to go beyond statistical tests on correlation and spread-stationarity, ensuring that the co-integration is supported by valid economic relationships (Hong and Hwang, 2021).
For each pair of stocks in our portfolio we first obtain an interaction feature vector $f_{i,j}$ by pairwise multiplying the feature vectors $s_{i},s_{j}$ representing the two company descriptions. Then we inspect the features that activate the most across all interaction vectors in our portfolio, and analyze them on the company descriptions of our most co-integrated pair using sae-vis (McDougall, 2024).
The co-integrated pair with the highest Sharpe ratio in $G_{\text{C-TM}}$ is UFP Technologies and Heico. Both share operational similarities in highly specialized manufacturing: UFP in custom-engineered products for medical and industrial applications; Heico produces aerospace, defense, and industrial components. Many interpretable and relevant features emerge, such as a ’renting’ feature 20155 highlighting both companies’ niche operations with logistical dependencies, including renting specialized facilities, exposing them to real estate fluctuations. We interpret the feature by examining its impact on the LLM’s logic in Appendix F.
3.4 Limitations
The following are the main limitations:
  1. 1.
    We do not fine-tune embedders, LLMs, nor SAEs. This is due to the high computational cost of supervised fine-tuning exceeding our computing budget, though fine-tuning could improve the performance of our method. This is an exciting direction for future work.
  2. 2.
    The exclusion of delisted stocks introduces survivorship bias into the analysis. This can be addressed with more complete data.
  3. 3.
    Our interpretability study is a proof of concept, limited by the availability of human feature annotators (there are tens of thousands of active features), and while we rely on the literature as support for the claim that SAE features are interpretable (Huben et al., 2024), our study could be expanded in the future.
4 Conclusions
We find that using SAE features is an effective and interpretable method for computing company similarity and explaining returns. Stress-testing, portfolio diversification, and other hedging-related strategies could all be subjects of further investigation using our clustering approach. We will release all our datasets and code in the camera ready.
Appendix A Clustering Embeddings
For BERT, we used bert-base-uncased from the transformers library. For SBERT, we used all-MiniLM-L6-v2 from the sentence_transformers library. For PaLM-gecko, we used textembedding-gecko@003 from the vertexai library.
Chunking: In our methodology, for both BERT and SBERT, we followed Vamvourellis et al. (2023) and implemented a chunking mechanism to accommodate the models’ maximum token limit of 512. Specifically, company descriptions exceeding this limit were split into overlapping chunks of 512 tokens. The [CLS] embeddings of these chunks were averaged to generate a single document embedding of 1536 tokens. For PaLM-Gecko, we leveraged its extended context window of 3072 tokens and directly processed the descriptions without chunking.
The pipeline below is optimised through Optuna’s Tree-structured Parzen Estimator (TPE) sampler for Bayesian hyperparameter optimization. The objective function maximizes $\text{MeanCorr}(G_{k})$. This search is constrained to 150 trials and a maximum timeout of 9 hours to balance thoroughness and resource usage:
(a) Dimensionality Reduction with UMAP: Given the high dimensionality of the input embeddings (768-dimensional vectors derived from a BERT model), we employ Uniform Manifold Approximation and Projection (UMAP) umap to reduce these high-dimensional textual embeddings to a lower-dimensional space, preserving both local and global data structures. We optimize three UMAP parameters to improve the quality of the downstream clustering: (1) n_components (target dimensionality); (2) n_neighbors; and (3) min_dist. All embeddings are standardized and casted to float32 to ensure computational efficiency.
(b) Clustering with Spectral Clustering: After reducing dimensionality, we perform clustering using Spectral Clustering, which is capable of handling noise and complex cluster shapes, following Vamvourellis et al. (2023). We first construct an affinity matrix from a k-nearest neighbors (KNN) graph of the UMAP outputs. Spectral Clustering then operates on this graph’s eigenstructure to form clusters. The number of clusters (n_clusters) is tuned via Optuna, while the neighborhood size ($k$) is set to a constant of 5, following Vamvourellis et al. (2023).
(c) Temporal Cross-Validation: To evaluate the stability and temporal generalization of the resulting clusters, we employ temporal cross-validation. The dataset is split into chronological folds. This setup reduces temporal bias and assesses whether the identified cluster structure remains consistent over time. We used parallel processing to evaluate each fold.
Embedder Cluster Group $(G_{\text{embedder}})$ UMAP $n_{\text{components}}$ UMAP $n_{\text{neighbors}}$ UMAP min_dist
$G_{\text{BERT}}$ 7 119 0.109
$G_{\text{SBERT}}$ 7 79 0.012
$G_{\text{PaLM-gecko}}$ 6 40 0.120
Table 2: Optimized UMAP Thresholds for Embedders
Embedder Cluster Group $(G_{\text{embedder}})$ Spectral $n_{\text{clusters}}$ Spectral $n_{\text{neighbors}}$
$G_{\text{BERT}}$ 10 5
$G_{\text{SBERT}}$ 49 5
$G_{\text{PaLM-gecko}}$ 27 5
Table 3: Optimized Spectral Clustering Thresholds for Embedders
Appendix B Clustering Sparse Features
We optimize the cutoffs using Optuna in cross validation. We observe that nearby nodes in the Minimum Spanning Tree (MST) are close in value, indicating a smooth distribution of distances. This is evident in the plot of the edge-weight cutoff threshold (x-axis) against $\text{MeanCorr}(G_{k})$ of the resulting cluster groups (y-axis), which shows a continuous and stable trend. The smoothness of the distribution suggests that the underlying sparse features are meaningfully grouped, and the cutoff thresholds effectively preserve local structure while removing noise. This property supports the reliability of the generated clusters and aligns with the ultrametric framework, ensuring interpretability and robustness.
Sparse Feature Cluster Group $(G_{\text{sparse\_feature}})$ MST edge-weight cutoff
$G_{\text{C-TS}}$ 0.1940
$G_{\text{C-TM}}$ 0.2315
Table 4: Optimized MST edge-weight cutoff Thresholds for Sparse Feature Cluster Group
Appendix C Proof of Manhattan metric
To show that $M_{i,j}$ satisfies positive definiteness, we verify the following:
1. Non-negativity: By definition,
$M_{i,j}=\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|\geq 0,$
as absolute differences are always non-negative.
Zero only for identical points: If $M_{i,j}=0$, then each term in the sum must be zero, i.e.,
$\text{PCA}_{k}(f_{i})=\text{PCA}_{k}(f_{j})\quad\forall k.$
Since the PCA transformation is invertible for features in its span, this implies $f_{i}=f_{j}$. Conversely, if $f_{i}=f_{j}$, then all differences are zero, so $M_{i,j}=0$.
2. Symmetry: Using the property of absolute values,
$\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|=\left|\text{PCA}_{k}(f_{j})-\text{PCA}_{k}(f_{i})\right|\quad\forall k.$
Therefore,
$M_{i,j}=M_{j,i}.$
3. Triangle Inequality:
For each dimension $k$, for all $l$, the absolute value satisfies the triangle inequality:
$\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|$ $\leq\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{l})\right|$
$\quad+\left|\text{PCA}_{k}(f_{l})-\text{PCA}_{k}(f_{j})\right|.$
Summing over all dimensions $k$ gives:
$\begin{array}[]{l}\displaystyle\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|\leq\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{l})\right|\\ \displaystyle\quad+\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{l})-\text{PCA}_{k}(f_{j})\right|.\end{array}$
This simplifies to:
$M_{i,j}\leq M_{i,l}+M_{l,j}.$
Thus, $M_{i,j}$ satisfies the triangle inequality.
Appendix D Trading details
In the out-of-sample evaluation (Jan 2014–Dec 2020), we simulated live trading of co-integrated pairs using a mean-reversion strategy.
Following Miao (2014), we assumed zero transaction costs, opening positions when the residual spread deviated beyond ±$\pm 1\sigma$ its mean, and closing when the spread reverted to the mean. A stop-loss mechanism is triggered if the spread exceeds ±$\pm 2\sigma$.
We allocated a fixed capital per pair and normalized the daily Total Portfolio Value (cash + unrealized PnL) to enable fair comparisons across clustering methods, regardless of the number of pairs. This normalization ensures that clusters with varying pair counts (e.g., 5 pairs vs. 60 pairs) have comparable portfolio value trajectories, where each pair is traded with an equivalent unit of capital (i.e., \$1). The daily normalized portfolio value series was subsequently used for Sharpe ratio calculations.
Appendix E Siamese Hyperparameters
Siamese Network was trained for 10 epochs with a learning rate of 0.001. The loss function used is Mean Squared Error (MSE). The hidden dimensions are 128 and 64 with a ReLu activation function.
Appendix F Interpretability of Feature 20155
Token Logit Contribution
tenant 0.28
tenants 0.24
Tenant 0.23
space 0.23
tenant 0.21
rents 0.21
shell 0.21
office 0.21
Table 5: Logit contribution on the top 8 tokens.